OpenStudy (lgbasallote):

another problem :/ Using the CYLINDRICAL SHELL METHOD find the volume generated by rotating the region bounded by \(x = 1 + (y - 2)^2\) and x = 2 about the x-axis. I got \(\frac{32\pi}{3}\) but the book says it's \(\frac{16\pi}{3}\)

OpenStudy (lgbasallote):

\[2\pi \int_1^3 y(5 + y^2 - 4y) dy\] this is the equation i used

OpenStudy (lgbasallote):

@apoorvk ? can you help mate?

OpenStudy (lgbasallote):

can you help me @SmoothMath this has been hanging out here long :/

OpenStudy (marco26):

\[2pi \int\limits_{1}^{3}y(-y^2+4y-3)dy\]

OpenStudy (lgbasallote):

why -3? @marco26 \(-2)^2 = 4

OpenStudy (lgbasallote):

\((-2)^2 = 4\) *

OpenStudy (lgbasallote):

unless you took out a negative...

OpenStudy (lgbasallote):

which seems not to be the case

OpenStudy (marco26):

Here's my equation:\[2pi \int\limits_{1}^{3}y[2-(y^2-4y+5)]\]

OpenStudy (lgbasallote):

where did you get 2 -(y^2 - 4y +5)?

OpenStudy (marco26):

the formula for horizontal strip for shell method is\[V=2pi \int\limits_{y1}^{y2}y_{c}(x_{R}-x_{L})dy\] where yc is the distance of the element from the axis of revolution. In your problem, Xr=2, and Xl=y^2+4y-5, and yc=y,, just plug them in

OpenStudy (marco26):

oops i mean Xl= y^2-4y+5

OpenStudy (lgbasallote):

what does r and l mean?

OpenStudy (lgbasallote):

ohhh this is the area formula right??

OpenStudy (marco26):

no, this is the formula for volume using shell method, given you have horizontal strip

OpenStudy (marco26):

X=Xr-Xl

OpenStudy (lgbasallote):

no..i meant xr - xl that's area bounded between two curves or something

OpenStudy (marco26):

here,XR-XL will be the height of the cylinder, yc is its radius, dy is the thickness

OpenStudy (lgbasallote):

uhmm you didnt really answer what i asked o.O

OpenStudy (anonymous):